Shift Summary for Eve Shift, 6/25/99

Sydney A. Taegar (taegar@physics.arizona.edu)
Fri, 25 Jun 1999 23:21:48 -0700 (MST)

Shift Summary for Eve shift, 6/25/99

On Shift: Albert Erwin, Sydney Taegar

Hours of beam delivered at all intensities: 7 hours
Hours of beam delivered at nominal intensity: 7 hours
Hours of E832 data (special runs not included): 5.5 hours
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time:

Beam studies, stopping and starting runs.

========================================================================
Runs ended this shift:
Run 13731 - 0.38M events - BEAM_832 (beam studies)
Run 13732 - 0.45M events - BEAM_832 (beam studies)
Run 13733 - 0.59M events - BEAM_832 (beam studies)
Run 13734 - 0.52M events - BEAM_832 (stopped to install new t0s for DCs)
Run 13735 - 7.0M events - BEAM_832 (stopped to update pipeline ped files)
Run 13736 - - BEAM_832 (in progress)

Tapes brought to FCC this shift: NONE

========================================================================
Accesses during shift: NONE

========================================================================
Overview of shift:

Rick Coleman's beam studies to check the alignment of the primary
collimator continued until 16:45. He found that the horizontal
alignment was fine but had to lower the primary collimator by
0.5 mm. See logbook XX.46-47.

The rest of the shift was standard BEAM_832 running.

RickK has modified EWUD logic to UD only until the MLU can
be replaced at the next downtime.

The 2V-L1 timing has shifted about +0.5 ns relative to nominal.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13731 15:30 RickC beam studies in progress.

15:35 RickK modified EWUD logic to UD-only until he
gets a chance to replace the MLU causing the
L1 inefficiency. This change will take effect
in run 13732. Details are in
/home/da_ktev/doc/vewud_99blues.txt and
logbook XX.48.

15:37 Stop run 13731.

13732 16:05 Start BEAM_832 run 13732 for RickC's beam studies.

16:20 Stop run 13732.

13733 16:30 Start BEAM_832 run 13733 for RickC's beam studies.
RickK remade L3 drivers prior to this run.

16:45 RickC is done with beam studies. Results
are summarized in shift overview above.
Stop run 13733.

13734 16:53 Start BEAM_832 run 13734.

17:12 Mounted tape QKA001 for onsplit.

17:25 Stop run 13734 so that Sasha Glazov can install
new t0s for the DCs to compensate for the
1ns change in RF timing yesterday.

13735 17:35 Start BEAM_832 run 13735.

16 No beam.

26 Beam is back.

18:15 VJ and LM note that EWUD problem might be caused
by inefficient channel V'(22) which overlaps with
both V0(7) and V0(23).

46 18:30 KTEVANA errors -> KFREND.

18:31 We notice that NM2SEED2 HMN = -12.9! Should be
-3.3 according to RickF, so the beam is ~1mm off!
Call MCR -- they are already on the case and
refer us to the Ch13 message:

"Left bends trouble. No estimate. Tuning
compensation for KTeV being made."

18:43 MCR calls - they are having trouble with AUTOTUNE
so they will turn it off and tune the beam manually
to correct the horizontal beam position.

62 18:51 NM2SEED2 HMN now -7.5.

80 19:15 NM2SEED2 HMN now -3.3 -- nominal position.

95 19:37 KTEVANA errors -> KFREND.

19:50 RickT shows up and notes factor of 2 variation
in beam intensity over a ~1sec period on the spill
scope. He also notes that the chamber ramping
is not critically damped.

20:50 We rediscover 35 MeV bump in CsI energy/block plot
mentioned in previous shift report. We also notice
that CsI channels 1642 and 2589 have high readout
occupancy. These two channels along with 1678 and
2310 had DPMTs swapped yesterday. We page
Ed -- he determines that the pipeline ped files were
not updated after the DPMT swap. We update the
files and stop run 13735.

13736 21:00 Start BEAM_832 run 13736 after brute force
pipeline initialization to load new ped files.

35 21:55 No beam.

44 22:06 Beam is back.

22:12 Beam structure still looks ratty. Call MCR -- they're
looking into it.

83 23:00 No beam.

23:05 Mounted tape QKSM02 for onsplit.

95 23:15 Beam is back and still ratty on the spill scope
although the BOA (~1.2) is fine. Call MCR --
they are still "looking into it." We should follow
up on this after the shift change-over.

23:40 From the DC2OR scaler display, Albert notices that
the 4th paddle in DCOR 2X seems to be dead. This
is most likely the scaler that is dead and should
be checked on the next access.

123 23:52 KTEVANA errors -> KFREND.

00:00 Hand off to owl shift with run 13736 in progress.

========================================================================
Follow up Notes on Previous Shift

RickK has changed EWUD logic to UD-only until he can replace
the MLU at the next downtime. This change took effect at run 13732.

35 MeV bump in CsI energy/block plot traced to incorrect pedestals
downloaded to pipeline for channels 1642 and 2589 (also 1678 and 2310
-- though these would show up as a deficit -- a bigger problem).
The DPMTs for these channels were swapped yesterday. The pipeline
pedestals have now been updated. This change took effect at run 13736.

========================================================================
Items to Follow up on Next Shift

o Take MUON run at next tape change.

o Follow up on beam structure with MCR.